Term Structure A 1-year Treasury bill currently offers a 4% rate of return. A 2-year Treasury note offers a 4.5% rate of return. Under the expectations theory, what rate of return do investors expect a 1-year Treasury bill to pay next year? 


The rate of return investors expect a 1-year Treasury bill to pay next year is _______ %. (Round to two decimal places.)
Answer #1

Let I% be the rate that is to be paid next year, then

(1+0.04) * (1+ i ) = (1+0.045)^2

(1.04) * (1+ i ) = (1.045)^2

1+i = 1.045^2 / 1.04 = 1.0500

I = 1.05 - 1 = 0.05 = 5%